Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured iReq 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B Compute (a) last years CM ratio and the break-even point in balls, and (b) the deReq 1 Req 2 Req3 Req 3 Req 4 Rega Reas Req 5 ReqGA Req GB Req 6A Req 6B Due to an increase in labor rates, the company estimaReq 1 Req 2 Req 3 Req 4 Rega Reg 5 Req5 Req 6A Req 6B Refer to the data in Required (2). If the expected change in variable eReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Refer again to the data in Required (2). The president feels that the company musRefer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plantReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B If the new plant is built, how many balls will have to be sold next year to earn| Rega Req3 Req 1 Req 2 Req 3 Req 4 Rega Reas Req 5 Req6A Req 6A | Req GB Req 6B IIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIII
Answer #1

C Amount А 1 Req 1(a) 2 Sales 3 Less: Variable expenses 4 Contribution Margin Per Unit 11,00,000.00 $ 25.00 6,60,000.00 $ 15.

Amount 1100000 660000 =B2-B3 Per Unit =B2/44000 =B3/44000 =B4/44000 Reg 4 Variable expenses CM Margin 792000 0.4 =F2/(1-F3) 1

Please do comment for any clarification. Thank you.